# Entanglement in quenched extended Su-Schrieffer-Heeger model with anomalous dynamical quantum phase transitions

### v1

@@ Appendix B Type-A anomalous DQPTs in the extended SSH model
Type-A anomalous DQPTs were first discovered back in 2017 on the study of quench and dynamical phase transitions in the transverse-field Ising model with long-range interactions, where a minimum in the Loschmidt rate function occurs sooner than the first critical time (the first occurrence of DQPT) [5, 72, 83, 84, 82, 6, 16, 85, 86]. In the extended SSH model, which is a topological model, we observe similar phenomenon for quenches between two different phases. Figure 17 shows the dynamical phase diagram for the type-A anomalous DQPTs in different initial phases, with the red regions indicating the regions where type-A anomalous DQPTs occur. First notice that there is no type-A anomalous DQPT happening for quenches from both $W_{2}$ and $W_{-1}$ phases, as seen from Fig. 17(c) and 17(d) where the red region is absent. In other words, type-A anomalous DQPTs only occur for quenches from class-1 ($W_{1},W_{0}$) phases (Fig. 17(a) and 17(b)).
@@
In Fig. 17(a) and 17(b), the type-B anomalous ($3k^{*}$) regions observed in Fig. 2(a) and 2(b) are outlined by the yellow lines. We find that a portion of the type-B anomalous DQPT phase boundary is shared with the type-A anomalous DQPT phase boundary. As one crosses this dynamical phase boundary from the $1k^{*}$ region to the $3k^{*}$ region, one of the two extra $k^{*}$’s can result in a critical time ($t^{*}_{B}$) much earlier than the original first critical time in the $1k^{*}$ region ($t^{*}_{A}$). As an example, Fig. 18(a) shows the LR for quenches from the $W_{1}$ phase to two final points indicated by the hollow triangles in Fig. 17(b). The red curve shows the case in the $1k^{*}$ region with type-A anomalous DQPT while the blue curve shows the case in the $3k^{*}$ region without type-A anomalous DQPT. The first smooth peak in the red curve might be considered as a pre-development of the nonanalytic peak at $t^{*}_{B}$, which arises from one of the two extra $k^{*}$’s when entering the $3k^{*}$ region, in the blue curve. If $t^{*}_{A}$ occurs much later than $t^{*}_{B}$, the LR may experience a minimum before reaching the nonanalyticity at $t^{*}_{A}$, and thus results in a type-A anomalous DQPT. However, the blue region where type-A anomalous DQPT is absent can also be found inside the $1k^{*}$ region, where the critical time is relatively small, as illustrated with the LR in Fig. 18(b).
One notices the concentration of the red regions around the $t_{d}=t_{c}$ equilibrium phase boundary in Fig. 17(a) and 17(b). Some selected quenches in these regions are demonstrated as red lines in Fig. 18(b) and 18(c). For the quench cases near this boundary, the critical times increase as the final points approach the boundary. Plus the closer the final points to the boundary, the more the number of local minima built up in the LRs, these DQPTs have to be type-A anomalous. In addition to this, the minimum achieved by these quenches in the extended SSH model is not the same as those studied in Ref. [5, 72, 83, 84, 6, 16]. Namely, the minima shown in the literature are in fact zeros of the LRs, while the minima presented in here have a finite value. The latter suggests the system at this time still deviates from the initial state even when the LR achieves a local minimum. This qualitative difference may be attributed to the topological characters of the present model as compared to a non-topological spin model. Further investigation into the correlation, if any, between the two types of anomalous DQPTs will be an interesting future work.
@@ Appendix C Entanglement echo rate function
Entanglement echo signals level crossings at the $\xi(t)=0.5$ Fermi level in the correlation matrix spectrum. It is defined as the overlap of the instantaneous entanglement ground state onto the initial entanglement ground state

### v2

@@ Appendix B Type-A anomalous DQPTs in the extended SSH model
Type-A anomalous DQPTs were first discovered back in 2017 on the study of quench and dynamical phase transitions in the transverse-field Ising model with long-range interactions, where a minimum in the Loschmidt rate function occurs sooner than the first critical time (the first occurrence of DQPT) [5, 72, 83, 84, 82, 6, 16, 85, 86, 100]. In the extended SSH model, which is a topological model, we observe similar phenomenon for quenches between two different phases. Figure 17 shows the dynamical phase diagram for the type-A anomalous DQPTs in different initial phases, with the red regions indicating the regions where type-A anomalous DQPTs occur. First notice that there is no type-A anomalous DQPT happening for quenches from both $W_{2}$ and $W_{-1}$ phases, as seen from Figs. 17(c) and 17(d) where the red region is absent. In other words, type-A anomalous DQPTs only occur for quenches from class-1 ($W_{1},W_{0}$) phases (Figs. 17(a) and 17(b)).
@@
In Figs. 17(a) and 17(b), the type-B anomalous ($3k^{*}$) regions observed in Figs. 2(a) and 2(b) are outlined by the yellow lines. We find that a portion of the type-B anomalous DQPT phase boundary is shared with the type-A anomalous DQPT phase boundary. As one crosses this dynamical phase boundary from the $1k^{*}$ region to the $3k^{*}$ region, one of the two extra $k^{*}$’s can result in a critical time ($t^{*}_{B}$) much earlier than the original first critical time in the $1k^{*}$ region ($t^{*}_{A}$). As an example, Fig. 18(a) shows the LR for quenches from the $W_{1}$ phase to two final points indicated by the hollow triangles in Fig. 17(b). The red curve shows the case in the $1k^{*}$ region with type-A anomalous DQPT while the blue curve shows the case in the $3k^{*}$ region without type-A anomalous DQPT. The first smooth peak in the red curve might be considered as a pre-development of the nonanalytic peak at $t^{*}_{B}$, which arises from one of the two extra $k^{*}$’s when entering the $3k^{*}$ region, in the blue curve. If $t^{*}_{A}$ occurs much later than $t^{*}_{B}$, the LR may experience a minimum before reaching the nonanalyticity at $t^{*}_{A}$, and thus results in a type-A anomalous DQPT. However, the blue region where type-A anomalous DQPT is absent can also be found inside the $1k^{*}$ region, where the critical time is relatively small, as illustrated with the LR in Fig. 18(b).
One notices the concentration of the red regions around the $t_{d}=t_{c}$ equilibrium phase boundary in Figs. 17(a) and 17(b). Some selected quenches in these regions are demonstrated as red lines in Figs. 18(b) and 18(c). For the quench cases near this boundary, the critical times increase as the final points approach the boundary. Plus the closer the final points to the boundary, the more the number of local minima are built up in the LRs, these DQPTs have to be type-A anomalous. In addition to this, the minimum achieved by these quenches in the extended SSH model is not the same as those studied in Ref. [5, 72, 83, 84, 82, 6, 16, 85, 86, 82, 85, 86, 100]. Namely, the minima shown in the literature are in fact zeros of the LRs, while the minima presented in here have a finite value. The latter suggests the system at this time still deviates from the initial state even when the LR achieves a local minimum. This qualitative difference may be attributed to the topological characters of the present model as compared to a non-topological spin model. Further investigation into the correlation, if any, between the two types of anomalous DQPTs, and the time evolution of the order parameters will be an interesting future work.
@@ Appendix C Entanglement echo rate function
Entanglement echo signals level crossings at the Fermi level ($\xi_{j}=0.5$) in the correlation matrix spectrum. It is defined as the overlap of the instantaneous entanglement ground state onto the initial entanglement ground state